OpenStudy (jjuden):

Which equation would best help solve the following problem: Gary kicks a field goal with an initial vertical velocity of 38 meters per second. How long will it take the football to hit the ground?

OpenStudy (jjuden):

a –16t^2+ 38t = 0 b 16t^2 + 38t = 0 c –4.9t^2 + 38t = 0 d –4.9t^2 – 38t = 0